# Fully distributed consensus control for stochastic multi-agent systems under undirected and directed topologies

### 14 Mar 2026 v1

@@ IV-B Fully distributed consensus control for stochastic MASs with the undirected topology
The proof is in Appendix A1. ∎
@@
$\begin{split}V_{3}(\vartheta_{t},t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (29)
@@
$\begin{split}dV_{3}&(\vartheta_{t},t)=\mathfrak{L}V_{3}(\vartheta_{t},t)dt+2\vartheta^{T}(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{3}(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V_{3}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)\\ &+2\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
@@
$\begin{split}\mathfrak{L}V_{3}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+2\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t)\\ &-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{2\lambda_{2}(\mathcal{L})}$ and $\gamma\geq\varepsilon$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (37) into $\mathfrak{L}V_{3}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{3}(\vartheta_{t},t)\\ \leq&\vartheta^{T}\!(t)[\mathcal{L}\otimes(A^{T}P+PA-2\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\!\vartheta^{T}\!(t)[\mathcal{L}\!\otimes\!(C^{T}\!PC)]\vartheta(t)\!-\!\gamma\sum_{i=1}^{N}\!e^{-\gamma t}(c_{i}(t)\!-\!\psi_{i})^{2}\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}\\ \leq&-\varepsilon V_{3}(\vartheta_{t},t).\end{split}$ (30)
@@
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V_{3}(\vartheta_{t},t)$, we have
@@
where $\delta>0$. Note that $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta_{t},t)$, where $\check{V}_{3}(\vartheta)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta$. Substituting the inequality $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta_{t},t)$ and (30) into the above equation yields
@@
$\begin{split}\mathrm{E}(e^{\delta t}\check{V}_{3}(\vartheta_{t}))\leq C_{3}.\end{split}$ (31)
That is, $\mathrm{E}(\check{V}_{3}(\vartheta_{t}))\leq C_{3}e^{-\delta t}$, which implies
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log\mathrm{E}(\check{V}_{3}(\vartheta_{t}))\leq-\delta.\end{split}$ (32)
@@
$\begin{split}&e^{\delta t}V_{3}(\vartheta_{t},t)-V_{3}(\vartheta_{0},0)\\ &=\int_{0}^{t}e^{\delta s}[\delta V_{3}(\vartheta(s),s)+\mathfrak{L}V_{3}(\vartheta_{s},s)]ds+M_{v3}(t),\end{split}$
@@
$\begin{split}e^{\delta t}\check{V}_{3}(\vartheta_{t})\leq C_{3}+M_{v3}(t),\end{split}$
Using the non-negative semi-martingale convergence theorem [38], ones have $\lim_{t\to\infty}\sup[e^{\delta t}\check{V}_{3}(\vartheta_{t})]<\infty$ a.s. Hence,
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log(\check{V}_{3}(\vartheta_{t}))\leq-\delta,a.s.\end{split}$ (34)
@@
Remark IV.6.
Theorem IV.4 proposed a fully distributed protocol for stochastic MASs to solve m.s. and a.s. consensus under undirected graphs. In comparison to the existing fully distributed stochastic consensus results[24, 26], we derive specific exponential estimates for the convergence rates of m.s. and a.s. consensus based on the proof of Theorem IV.4, that is, the relationship between the exponential convergence rate and system parameters is established.
Remark IV.7.
The consensus algorithm for undirected graphs proposed in this work is inspired by the protocol presented in [25]. While both address the consensus problem under similar stochastic settings, our contributions extend beyond the existing works, establishing a more comprehensive theoretical foundation:
1) Theoretical completeness: In contrast to [25], which primarily focuses on algorithm design and convergence analysis, we provide a rigorous proof of existence and uniqueness of solutions for the closed-loop stochastic differential system. This is accomplished by establishing verifiable sufficient conditions (Lemma III.1) within a stochastic Lyapunov framework.
2) Practical feasibility: We establish the uniform boundedness of the proposed protocol under the parameter constraint $\gamma<\frac{3}{2\lambda_{\max}(P)}$. This result guarantees that the control inputs remain bounded throughout the evolution of the system, a critical property for practical implementation that was not explicitly analyzed in [25].
3) Topological generality: The core theoretical contribution lies in the development of a fully distributed consensus framework applicable to directed graphs (Theorems IV.1 and IV.2). By incorporating the auxiliary time-varying gain function $\Sigma_{i}(t)$, our design eliminates the reliance on global network information and the assumption of topological symmetry, which is typically required for undirected graph cases like [25]. This extension significantly broadens the applicability to realistic networked systems with asymmetric information exchange.
Consequently, while inspired by [25], this work not only consolidates the theoretical basis for undirected topologies but also provides a novel and practical fully distributed approach for the more general directed case in stochastic environments.
Remark IV.8.
The protocol discussed in this section is a node-based adaptive protocol. Compared to edge-based adaptive protocols, this approach generally exhibits lower dimensionality since the number of edges in a connected graph typically exceeds that of its nodes. However, it is worth noting that edge-based adaptive protocols possess their own advantages, such as being more applicable to switching topology because the selection of Lyapunov function does not rely on the Laplacian matrix of a fixed topology. This aspect represents a potential avenue for future research as indicated in [24].
Remark IV.9.
The framework established in this paper can be naturally extended to nonlinear systems. While the current analysis focuses on linear dynamics with path-dependent adaptive feedback, the core analytical approach, including the construction of Lyapunov functions, the application of functional $\mathrm{It\hat{o}}$ formula, and the use of existence and uniqueness, essentially does not rely on the linear dynamics. For nonlinear dynamics satisfying appropriate growth conditions (such as, local Lipschitz condition and polynomial growth), a similar fully distributed protocol design and consensus analysis can be carried out for this class of stochastic nonlinear MASs. Such extensions will be discussed in future work.
@@ V-A The case of directed topology
Figure 3: The adaptive gains
Figure 4: The relative states errors and m.s. relative state errors
@@ V-B The case of undirected topology
Figure 5: The adaptive gains
Figure 6: The relative states errors and m.s. relative state errors
@@
Figure 7: The m.s. relative state errors for the case of $\gamma=0,0.5,1$
@@ VI Conclusion
The design of a unified fully distributed consensus protocol for a class of stochastic MASs under directed and undirected graphs are investigated. The existence and uniqueness of solutions for path-dependent and highly nonlinear closed-loop stochastic MASs under fully distributed protocols are first explored. For the case of directed graphs, a unified fully distributed control protocol is designed for the first time to solve m.s. and a.s. consensus for stochastic MASs under directed graphs. Then, for the case of undirected graphs, we develop an enhanced fully distributed protocol and derive explicit exponential convergence rates for both m.s. and a.s. consensus, precisely characterizing their dependence on system parameters. Future research directions include extending these results to MASs with compound disturbances and investigating more challenging scenarios involving heterogeneous nonlinear dynamics and time-varying communication topologies.
A1: The proof of Theorem IV.3
Proof.
Then, we choose Lyapunov function
$\begin{split}V_{2}(\vartheta_{t},t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\frac{1}{2}\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (36)
$\begin{split}dV_{2}&(\vartheta_{t},t)\!=\!\mathfrak{L}V_{2}(\vartheta_{t},t)dt+2\vartheta^{T}\!(t)(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{2}(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V_{2}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)+(c_{i}(t)-\psi_{i})\dot{c}_{i}(t).\end{split}$
Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Then, we can obtain
$\begin{split}\mathfrak{L}V_{2}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t).\end{split}$
Note that $\xi_{i}(t)=\sum_{j\in N_{i}}(x_{i}(t)-x_{j}(t))=\sum_{j\in N_{i}}(\vartheta_{i}(t)-\vartheta_{j}(t))$. Then, we have
$\begin{split}&\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t)\\ &=\vartheta^{T}(t)[\mathcal{L}(\tilde{C}(t)-\tilde{\psi})\mathcal{L}\otimes PBB^{T}P]\vartheta(t),\end{split}$ (37)
where $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{\lambda_{2}(\mathcal{L})}$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (37) into $\mathfrak{L}V_{2}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{2}(\vartheta_{t},t)\\ \leq&\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA-\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\vartheta^{T}(t)[\mathcal{L}\otimes(C^{T}PC)]\vartheta(t)\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)\\ \leq&-\varepsilon V_{2}(\vartheta_{t},t)+\bar{\varphi}(t),\end{split}$ (38)

### 31 Mar 2026 v2

@@ IV-B Fully distributed consensus control for stochastic MASs with the undirected topology
Then, we choose Lyapunov function
$\begin{split}V_{2}(\vartheta(t),t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\frac{1}{2}\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (27)
$\begin{split}dV_{2}&(\vartheta(t),t)\!=\!\mathfrak{L}V_{2}(\vartheta(t),t)dt+2\vartheta^{T}\!(t)(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{2}(\vartheta(t),t)$ is defined as
$\begin{split}\mathfrak{L}V_{2}&(\vartheta(t),t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)+(c_{i}(t)-\psi_{i})\dot{c}_{i}(t).\end{split}$
Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Then, we can obtain
$\begin{split}\mathfrak{L}V_{2}&(\vartheta(t),t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\vartheta_{i}^{T}(t)PBB^{T}P\vartheta_{i}(t).\end{split}$
Note that $\vartheta_{i}(t)=\sum_{j\in N_{i}}(x_{i}(t)-x_{j}(t))=\sum_{j\in N_{i}}(\vartheta_{i}(t)-\vartheta_{j}(t))$. Then, we have
$\begin{split}&\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\vartheta_{i}^{T}(t)PBB^{T}P\vartheta_{i}(t)\\ &=\vartheta^{T}(t)[\mathcal{L}(\tilde{C}(t)-\tilde{\psi})\mathcal{L}\otimes PBB^{T}P]\vartheta(t),\end{split}$ (28)
where $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{\lambda_{2}(\mathcal{L})}$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (28) into $\mathfrak{L}V_{2}(\vartheta(t),t)$ yields
$\begin{split}\mathfrak{L}&V_{2}(\vartheta(t),t)\\ \leq&\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA-\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\vartheta^{T}(t)[\mathcal{L}\otimes(C^{T}PC)]\vartheta(t)\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)\\ \leq&-\varepsilon V_{2}(\vartheta(t),t)+\bar{\varphi}(t),\end{split}$ (29)
@@
$\begin{split}V_{3}(\vartheta(t),t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (33)
@@
$\begin{split}dV_{3}&(\vartheta(t),t)=\mathfrak{L}V_{3}(\vartheta(t),t)dt+2\vartheta^{T}(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{3}(\vartheta(t),t)$ is defined as
$\begin{split}\mathfrak{L}V_{3}&(\vartheta(t),t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)\\ &+2\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
@@
$\begin{split}\mathfrak{L}V_{3}&(\vartheta(t),t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+2\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\vartheta_{i}^{T}(t)PBB^{T}P\vartheta_{i}(t)\\ &-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{2\lambda_{2}(\mathcal{L})}$ and $\gamma\geq\varepsilon$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (28) into $\mathfrak{L}V_{3}(\vartheta(t),t)$ yields
$\begin{split}\mathfrak{L}&V_{3}(\vartheta(t),t)\\ \leq&\vartheta^{T}\!(t)[\mathcal{L}\otimes(A^{T}P+PA-2\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\!\vartheta^{T}\!(t)[\mathcal{L}\!\otimes\!(C^{T}\!PC)]\vartheta(t)\!-\!\gamma\sum_{i=1}^{N}\!e^{-\gamma t}(c_{i}(t)\!-\!\psi_{i})^{2}\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}\\ \leq&-\varepsilon V_{3}(\vartheta(t),t).\end{split}$ (34)
@@
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V_{3}(\vartheta(t),t)$, we have
@@
where $\delta>0$. Note that $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta(t),t)$, where $\check{V}_{3}(\vartheta)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta$. Substituting the inequality $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta(t),t)$ and (34) into the above equation yields
@@
$\begin{split}\mathrm{E}(e^{\delta t}\check{V}_{3}(\vartheta(t)))\leq C_{3}.\end{split}$ (35)
That is, $\mathrm{E}(\check{V}_{3}(\vartheta(t)))\leq C_{3}e^{-\delta t}$, which implies
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log\mathrm{E}(\check{V}_{3}(\vartheta(t)))\leq-\delta.\end{split}$ (36)
@@
$\begin{split}&e^{\delta t}V_{3}(\vartheta(t),t)-V_{3}(\vartheta_{0},0)\\ &=\int_{0}^{t}e^{\delta s}[\delta V_{3}(\vartheta(s),s)+\mathfrak{L}V_{3}(\vartheta_{s},s)]ds+M_{v3}(t),\end{split}$
@@
$\begin{split}e^{\delta t}\check{V}_{3}(\vartheta(t))\leq C_{3}+M_{v3}(t),\end{split}$
Using the non-negative semi-martingale convergence theorem [38], ones have $\lim_{t\to\infty}\sup[e^{\delta t}\check{V}_{3}(\vartheta(t))]<\infty$ a.s. Hence,
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log(\check{V}_{3}(\vartheta(t)))\leq-\delta,a.s.\end{split}$ (38)
@@
While the design of the fully distributed protocol for undirected graphs in Theorem IV.4 is inspired by [25], our work differs from the results of [25] in the following three aspects. First, we provide a rigorous proof of the existence and uniqueness of solutions for the path-dependent stochastic system by establishing verifiable sufficient conditions (Lemma III.1) within a stochastic Lyapunov framework, which serves as the foundation for the fully distributed protocol design. Second, we establish the uniform boundedness of the proposed protocol under the parameter constraint $\gamma<\frac{3}{2\lambda_{\max}(P)}$, ensuring that the control inputs remain bounded throughout system evolution, a critical practical property not explicitly analyzed in [25]. Third, our core contribution lies in developing a fully distributed consensus framework for directed graphs (Theorems IV.1 and IV.2). By incorporating the auxiliary time-varying gain $\Sigma_{i}(t)$, the proposed protocol eliminates reliance on global network information and topological symmetry, thereby extending applicability to realistic networked systems with asymmetric information exchange.
@@ V-A The case of directed topology
Figure 3: The adaptive gains for directed topology
Figure 4: The adaptive gains for undirected topology
@@ V-B The case of undirected topology
Figure 5: The relative states errors and m.s. relative state errors for directed topology
Figure 6: The relative states errors and m.s. relative state errors for undirected topology
@@
Figure 7: The m.s. relative state errors $E\|x_{i1}(t)-x_{11}(t)\|^{2}$} for the case of $\gamma=0,0.5,1$
Figure 8: The m.s. relative state errors $E\|x_{i2}(t)-x_{12}(t)\|^{2}$} for the case of $\gamma=0,0.5,1$
@@ VI Conclusion
The design of a unified fully distributed consensus protocol for a class of stochastic MASs under directed and undirected graphs is investigated. The existence and uniqueness of solutions to path-dependent and highly nonlinear stochastic systems are first explored. For the case of directed graphs, a unified fully distributed control protocol is designed for the first time to solve m.s. and a.s. consensus for stochastic MASs. Then, for the case of undirected graphs, we develop an enhanced fully distributed protocol and derive explicit exponential convergence rates for both m.s. and a.s. consensus.
Future research directions include extending these results to MASs with compound disturbances and investigating more challenging scenarios involving heterogeneous nonlinear dynamics and time-varying communication topologies. Notably, the framework established in this paper can be naturally extended to nonlinear systems. While the current analysis focuses on linear dynamics with path-dependent adaptive feedback, the core analytical approach does not rely on the linear dynamics. For nonlinear dynamics satisfying appropriate growth conditions (such as, local Lipschitz condition and polynomial growth), a similar fully distributed protocol design and consensus analysis can be carried out for this class of stochastic nonlinear MASs. These extensions will be explored in future work.